A car starts from rest and reaches to a top speed of 15m/s in 15 seconds.An acceleration increases uniformly for the first 6 seconds and after that it remains constant. Compute the distance travelled in 15 seconds.?

Grade:12

1 Answers

Arun
6 years ago
Dear Ram
 
Let the acc = a for 6 sec.
Hence
V6 = 0 + a*6
V6 = 6a
Now
 
V15 = V6 + 0*9
V15 = 15 m/sec= V6
6a = 15
a = 2.5 m/s²
Hence
Distance travelled 
S = 0*6 + (1/2)*2.5* 6²
S = 45 m
 
Regards
Arun (askIITians forum expert)